OpenStudy (anonymous):

Find the equation of the tangent line to the curve y = 3tan x at the point (pi/4 , 3). The equation of this tangent line can be written in the form y = mx+b where m = b =

OpenStudy (anonymous):

First we want to find the derivative of y = 3tanx. You know how to do that?

OpenStudy (anonymous):

dy/dx=2sec

OpenStudy (anonymous):

No. If y = 3tanx Then y' = 3 sec^2(x)

OpenStudy (anonymous):

The first derivative gives us the slope. So use it to find the slope our tangent line will need to have at the point (pi/4, 3)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I need m (the slope and b(yaxis)?

OpenStudy (anonymous):

yes. first find m..the slope. then plug in m and (pi/4, 3) into y=mx + b and solve for b.

OpenStudy (anonymous):

m=3

OpenStudy (anonymous):

b=3

OpenStudy (anonymous):

y=3(pi/4)+3

OpenStudy (anonymous):

3=3(pi/4)+B

OpenStudy (anonymous):

need more help please